# Voting by mail: a Markov chain model for managing the security risks of absentee voting

### v1

@@ Abstract
The scrutiny surrounding vote-by-mail (VBM) in the United States has increased in recent years, raising concerns about the integrity and security of absentee voting. This paper addresses these issues by introducing a dynamic mathematical modeling framework for performing a risk assessment of VBM processes. We introduce a discrete-time Markov chain (DTMC) to model the VBM process and assess election performance and risk with a novel layered network approach that considers the interplay between VBM processes, malicious and non-malicious threats, and security mitigations. The time-inhomogeneous DTMC framework captures dynamic risks and evaluates performance over time. The DTMC model accounts for a spectrum of outcomes, from unintended voter errors to sophisticated, targeted attacks, representing a significant advancement in the risk assessment of VBM planning and protection. A case study based on real-world data from Milwaukee County, Wisconsin, is used to evaluate the DTMC model. The analysis includes the development of attack scenarios to assess the system’s resilience and the evaluation of security measures. The analysis suggests that ballot drop boxes and automatic ballot notification systems are crucial for ensuring secure and reliable operations.
@@ 1 Introduction
Democracies rely on the ability of the general public to vote on critical issues pressing their society. In the United States (US), there are generally two methods by which the public can cast a ballot. Voters can travel to their designated polling location where they receive and cast a ballot on an election day, referred to as in-person voting. Voters can also vote absentee, and many of those who vote absentee choose to accept and cast a ballot through the United States Postal Service (USPS), the vote-by-mail (VBM) system. Historically, in-person voting has been the primary method of voting in the US. However, VBM has experienced significant growth in recent years, accounting for 20-percent of votes cast in the 2016 General Election in the US and 46-percent of votes cast in the 2020 General Election that occurred during the COVID-19 pandemic [6].
@@
This paper presents a methodology to support VBM risk analysis by introducing a novel discrete-time Markov chain (DTMC) for VBM to capture dynamic election performance. The modeling approach employs attack tree models of malicious (intentional) and non-malicious (unintentional) risks and security mitigations. Attack tree models have been used to characterize risks in prior studies on voting by mail security [32, 23]. A novel aspect of the DTMC framework is that it models the linkages between attack trees, security mitigations, and VBM processes to characterize and predict dynamic VBM performance under various threat scenarios and operating conditions. Furthermore, the DTMC model leverages attack trees that provide a structured approach to capturing malicious and non-malicious risks to VBM systems. The DTMC framework highlights how different types, intentions, and timings of attacks impact overall VBM performance and security. Using real-world data from Milwaukee County in the 2020 General Election, we calibrate and validate the model and demonstrate its applicability to forecast the performance of the VBM process under different conditions and policy scenarios. Contributions of this paper include the following.
@@ 2 Literature review
Attack and fault trees have been used for the risk analysis of in-person voting processes. Yasinisac and Pardue [40] illustrate how attack trees have been deployed to analyze risks, including malicious cyber-attacks targeting voting machines. These methodologies consider both equipment failures and cybersecurity vulnerabilities, offering a broad view of potential threats. In a similar vein, Simidchieva et al. [28] use fault trees to evaluate the in-person election process, with a particular emphasis on ballot counting and vulnerability detection. However, these papers do not differentiate between malicious and non-malicious attacks, solely focus on in-person voting, and represent risks as static by excluding temporal analysis.
@@ 3.1 Process Layer
First, we introduce seven process states that capture the movement of ballots in the process layer and define the operation of the VBM system. Each of these process states is also a state in the DTMC, and the states represent the change in physical location of the ballot over time. The voter requests a ballot in state $I$, the election office then fulfills this request in state $II$ and mails and unmarked ballot to the voter. In state $III$ the unmarked ballot is handled by the United States Postal Service (USPS) and is in transit to the voter. In state $IV$ the unmarked ballot reaches the voter, however, it is up to the voter when the marked ballot is filled out and returned. In state $V$ the voter selects the method of return for the marked ballot. Standard operation for returning a ballot is a transition to state $VI$ via the USPS. Alternatively, the voter could return the ballot by drop box, if drop boxes are available. The final process state $VII$ occurs when the marked ballot reaches the election office and is held until processing on election day. The ballot process is summarized by the following states:
@@ 3.2 Final Ballot States
Note that the preferred ballot status is for ballots to reach the (C,U) state.
@@ 3.3 Attacks Layer
The attacks layer is comprised of active vulnerabilities to the VBM process. Vulnerabilities in the VBM system are documented in attack trees by the US Election Assistance Commission [32] and have been expanded upon since the 2020 General Election [22]. The terminal nodes of these attack trees represent “access” points to additional vulnerabilities within the process. For this reason, we model each of these terminal leaf nodes of the attack trees. Attack trees, a valuable tool in cybersecurity, help assess threats to a system [26]. They depict potential vulnerabilities and attacker paths in a Boolean logic tree structure, where the root node represents the ultimate goal of an attacker. Attack trees use a combination of Boolean AND($\times$) and OR($+$) logic gates to trace paths from terminal attack nodes back to the root node. Figure 1 shows an example of an attack tree that visualizes two attacks corresponding to the malicious loss ($X{13}$) and accidental loss ($X{14}$) of a ballot. In Figure 1, terminal attack nodes $X{13}$ and $X{14}$ form part of an attack path leading to a successful insider attack, represented by root node 1 with only OR-gates along the paths. These attack paths offer a static view of threats within the VBM process, with the terminal leaf nodes $X{13}$ and $X{14}$ serving as entry points.
@@ 3.4 Mitigations Layer
Label Mitigation description Controlling entity
M3 The ability of a voter to attain the status of their ballot. This has a high probability if automatic notifications are provided via BalloTrax/Ballot Scout. Election office
M4 Replacement ballot package request Voter
M5 Notify voter to send ballot back before deadline Election office
M6 In-person absentee voting Voter
M7 Return ballot via drop box Voter
This paper necessitates a thorough understanding of the connections between terminal node attacks in the VBM process and the available mitigation strategies designed to counteract them. To facilitate this, we employ a framework that examines the impact of terminal node attacks on the VBM process and incorporates recourse actions available to counter these attacks. Table 3 illustrates the relationship between attacks and mitigations. In this table, an “O” indicates that a mitigation strategy can counter the threat without impacting the ballot processing time. In contrast, a “D” signifies that while mitigation can counter the threat, it introduces a delay in the VBM process. For example, the scenario of a lost ballot ($X{13}$), a type of malicious attack, causes a delay resulting from the mitigation logic that includes the time required to notify the voter about their missing ballot followed by the time it takes for the voter to decide on an appropriate recourse action, such as M4 or M6. This scenario exemplifies how Table 3 summarizes key linkages in the DTMC model introduced in the next section. Consistent with the attacks in Table 1, terminal attack nodes listed in bold are incorporated into the DTMC model and are included in the computational results.
@@ 4 DTMC Model
Figure 2: Layered Network for time intervals $11\leq t<T-$ 1
@@ 5 Wisconsin 2020 General Election Case Study
We conduct a case study based on data from the 2020 General Election in Milwaukee County, Wisconsin. In 2020 Milwaukee County, the largest county in Wisconsin, had 939,489 residents and 478 voting wards [31, 4, 14]. Milwaukee County election officials used multiple absentee voting mechanisms during the 2020 General Election, including VBM with designated drop boxes and in-person absentee voting. The state of Wisconsin allowed absentee voting with “no excuses” and sent voter reminders to return ballots [16]. At that time, the county had 550,132 registered voters and experienced a turnout of 83.67 percent [36].
@@ 5.2 Transition probabilities
Table 4 outlines the transition probabilities for each arc active in the baseline model. The first and second columns define the two DTMC states associated with a non-zero transition probability. The next column provides a brief functional description of the ballot state transition that the arc represents. The columns labeled Interval 1, 2 and 3 represent the values of associated transition probability across time intervals 1, 2, and 3. We omit interval 4, since these transition probabilities are 00 or $1$ with arcs illustrated in Figure 3. The justification column reports a reference for each transition probability value and denotes which arcs were adjusted through calibration. The arc configuration aligns with the model shown in Figure 2.
We discuss several transition probabilities. The transition probability from process state $III$ to $IV$ is $0.938$, which reflects the proportion of election mail that is processed on-time as reported by the USPS Inspector General [19]. The rest of the ballots remain in state $III$. The transition from state $IV$ to $V$ reflects the return of completed ballots by voters (that is, the voter rate of ballot return). This transition captures the proportion of voters who return the marked ballot within a day of receiving it in the mail, which is determined as a part of the model calibration since this value is not directly recorded. MIT Election Data + Science Lab [15] provide the values for utilization of drop boxes over USPS to return ballots. Transitions from state $V$ to state $VII$, which occur with a probability of $0.515$, represents the proportion of ballots returned by drop boxes. [15] Transitions at the end of the time interval move ballots to their final ballot states with probability 1.0. These probabilities can be gleaned from Figure 3. All other transition probabilities are zero. We calibrated the DTMC using historical data from the WEC 2020 General Election report [38]. This involved adjusting the voter rate of ballot return, the arc probability from node $IV$ to $V$, to align the output of (C,U) and (NC,U) ballots with the observed ballot return rate seen in aggregate by the state of Wisconsin. Following calibration, we validate the model’s predictive accuracy by testing it against independent data validation points for Milwaukee County. This ensures blind prediction input data such that our model fits historical data and ensures the model is capable of making accurate predictions in varying scenarios.
@@ 5.3 Calibration
Mitigation
Mitigation strength Justification
@@
M3 0.0265 status were not implemented
@@
M4 0.900 replacement ballots
M5 0.740 Region used ballot return reminders
M6 0.400 In-person absentee was implemented
@@
M7 0.520 Elections Dataverse [3]
@@
The calibration of the baseline DTMC model also requires consideration of non-malicious attacks; we model as part of normal system operations. The ballot rejection rate of the model is associated with the arc transition probability from node $IV$ to $(X{65}+X{67}+X{68})$. It correlates directly with the number of ballots rejected by the election office due to voter error. We set the ballot rejection rate to 0.000162 to closely match the observed number of returned ballots that were unaltered and not counted (NC,U). Next, we vary the impact of the remaining non-malicious attacks, representing the ballots that never returned to the election office. Recall terminal attack node $X14$ represents the accidental loss of ballots in the mail. The terminal attack node $X30$ is the misaddressing of ballots to the voter, and $X36$ represents the accidental loss of ballots in the election office. These non-malicious attacks all result in the same penalty, and the ballots impacted are considered not counted and not returned (NC,NR). Unlike the ballot rejection rate of the election office, there is a lack of evidence to support the different magnitudes for the non-malicious attacks $X{14},X{30},X{36}$. These non-malicious attack strength values are equivalent for all periods $t=1,...,T-1$ [35]. Therefore, we set the terminal attacks $X{14},X{30},X{36}$ to equal strength. As a result, the following arcs have the same transition probability of 0.0343: arc from $IV$ to $X14$; arc from $II$ to $X30$; arc from $VII$ to $X36$. These arcs result in a mean absolute deviation of 0.14-percent from the observed (C,U) ballots.
@@ 5.4 Validation
v5 All sent ballots County specific sum of all ballot states 2,059,130 325,547 325,351
@@ 6 Computational Results
This section discusses the baseline DTMC model results for the Milwaukee County cawe study. We then study the impact of several types of malicious attacks on the established baseline. We report the expected number of ballots in each of the final ballot states—rounded to the nearest ballot—under multiple scenarios.
@@ 6.1 Baseline model
We first analyze the VBM system under normal conditions with non-malicious attacks in the baseline model introduced in Section 5. Non-malicious attacks (see Table 4) are included in the baseline model, since they are an uncontrolled, continuous part of the VBM process. Additionally, we perform a sensitivity analysis of four of the non-malicious attacks by varying their corresponding probabilities of occurring.
@@
Election office $X{14},X{30},X{36}$ Final Ballot States (Expected number of ballots) Scenario rejection rate attack strength (C,U) (NC,U) (C,A) (NC,L) (NC,NR) (NC,A) Baseline 0.000162 0.0343 324,792 559 0 173 20,371 0 Baseline that varies election office rejection rate 0.0000707 0.0343 325,091 245 0 174 20,386 0 from node $IV$ to $(X{65}+X{67}+X{68})$ 0.000111 0.0343 324,958 385 0 174 20,379 0 0.000172 0.0343 324,759 594 0 173 20,369 0 0.000424 0.0343 323,933 1,464 0 173 20,325 0 0.000818 0.0343 322,653 2,812 0 172 20,258 0 Baseline that varies non-malicious attack strength 0.000162 0.00202 339,228 559 0 198 5,911 0 associated with $X14,X30,X36$ 0.000162 0.0192 331,493 559 0 184 13,660 0 0.000162 0.0384 323,025 559 0 171 22,141 0 0.000162 0.0576 314,732 560 0 158 30,445 0 0.000162 0.0939 299,469 560 0 140 45,728 0
The top row of Table 7 represents the baseline model, and it reports the expected number of ballots in each final ballot state. Each subsequent row represents a different scenario with varying attack parameters. The incremental increase of the rejection rate associated with the arc transition probability from node $IV$ to $(X{65}+X{67}+X{68})$ by a probability of 0.00001 increases the number of (NC,U) ballots by 6-percent (35 total ballots), suggesting that efforts to reduce the rejection rate can decrease the number of ballots that are not counted. Higher attack strengths associated with $X{14},X{30},X{36}$ increase the number of ballots not returned (NC,NR). For example, when the attach strength increases from $0.0343$ to $0.0576$, the number of (NC,NR) ballots increase by $10,074$. Additionally, we note that the election office rejection rate ($X65+X66+X67$) for Milwaukee County is substantially lower than that of the other non-malicious attacks. These insights underscore the need for additional attention to unreturned (NC,NR) ballots in the VBM process and maintaining minimal non-malicious attack strengths to ensurethat ballots are returned and counted.
@@ 6.2 Malicious attack scenarios given baseline
Next, we consider introduce malicious attacks to the baseline. We focus on malicious attacks $X9,X13,$ and $X29$. Malicious attacks are modeled to last a single day during the 49-day time horizon to reflect feasible real-world election interference events, and we vary the attack strength of each attack to consider different scenarios. Malicious attacks therefore, modify the transition probability matrix ($P_{t^{*}}$) for a single day associated with the attack $t^{*}$. This approach allows us to assess the impact of attack timings and the mitigation strategies for countering attacks.

### v2

@@ Abstract
The scrutiny surrounding vote-by-mail (VBM) in the United States has increased in recent years, raising concerns about the integrity and security of absentee voting. This paper addresses these issues by introducing a dynamic mathematical modeling framework for performing a risk assessment of VBM processes. We introduce a discrete-time Markov chain (DTMC) to model the VBM process and assess election performance and risk with a novel layered network approach that considers the interplay between VBM processes, malicious and non-malicious threats, and security mitigations. The time-inhomogeneous DTMC framework captures dynamic risks and evaluates performance over time. The DTMC model accounts for a spectrum of outcomes, from unintended voter errors to sophisticated, targeted attacks, representing a significant advancement in the risk assessment of VBM planning and protection. A case study based on real-world data from Milwaukee County, Wisconsin, is used to evaluate the DTMC model. The analysis includes the development of attack scenarios and the evaluation of security measures, to illustrate the impact of different attack timings. The analysis suggests that ballot drop boxes and automatic ballot notification systems are crucial for ensuring secure and reliable operations.
@@ 1 Introduction
The legitimacy of any political system hinges on the accuracy and fairness of its voting processes. In the United States (US), there are generally two methods by which the public can cast a ballot. Voters can travel to their designated polling location where they receive and cast a ballot on an election day, referred to as in-person voting. Voters can also vote absentee, and many of those who vote absentee choose to accept and cast a ballot through the United States Postal Service (USPS), the vote-by-mail (VBM) system. Historically, in-person voting has been the primary method of voting in the US. However, VBM has experienced significant growth in recent years, accounting for 20-percent of votes cast in the 2016 General Election in the US and 46-percent of votes cast in the 2020 General Election that occurred during the COVID-19 pandemic [6].
@@
This paper presents a methodology to support VBM risk analysis by introducing a novel discrete-time Markov chain (DTMC) for VBM to capture dynamic election performance. The modeling approach uses the attack tree models of malicious (intentional) and non-malicious (unintentional) risks and security mitigations. Attack tree models have been used to characterize risks in prior studies on voting by mail security [34, 24]. A novel aspect of the DTMC framework is that it models the linkages between attack trees, security mitigations, and VBM processes to characterize and predict dynamic VBM performance under various threat scenarios and operating conditions. Furthermore, the DTMC model leverages attack trees that provide a structured approach to capturing malicious and non-malicious risks to VBM systems. We focus on the number of ballots affected by election risks, not on the likelihood of the outcome of an election changing, since any inaccuracies weaken the legitimacy of a political system, regardless of whether these inaccuracies change election outcomes. The DTMC framework highlights how different types, intentions, and timings of attacks impact overall VBM performance and security. We acknowledge that successful attacks could change the outcome of an election. Using real-world data from Milwaukee County in the 2020 General Election, we evaluate VBM performance under different conditions and policy scenarios. Contributions of this paper include the following.
@@ 2 Literature review
Attack and fault trees have been used for the risk analysis of in-person voting processes. Yasinsac and Pardue [42] illustrate how attack trees have been deployed to analyze risks, including malicious cyber-attacks targeting voting machines. These methodologies consider both equipment failures and cybersecurity vulnerabilities, offering a broad view of potential threats. In a similar vein, Simidchieva et al. [29] use fault trees to evaluate the in-person election process, with a particular emphasis on ballot counting and vulnerability detection. However, these papers do not differentiate between malicious and non-malicious attacks, solely focus on in-person voting, and represent risks as static by excluding temporal analysis.
@@ 3.1 Process Layer
First, we introduce seven process states that capture the movement of ballots in the process layer and define the operation of the VBM system. Each of these process states is also a state in the DTMC, and the states represent the change in physical location of the ballot over time. The voter requests a ballot in state $I$, the election office then fulfills this request in state $II$ and mails an unmarked ballot to the voter. In state $III$ the unmarked ballot is handled by the United States Postal Service (USPS) and is in transit to the voter. In state $IV$ the unmarked ballot reaches the voter, however, it is up to the voter when the marked ballot is filled out and returned. In state $V$ the voter selects the method of return for the marked ballot. Standard operation for returning a ballot is a transition to state $VI$ via the USPS. Alternatively, the voter could return the ballot by drop box, if drop boxes are available. The final process state $VII$ occurs when the marked ballot reaches the election office and is held until processing on election day. The ballot process is summarized by the following states:
@@ 3.2 Final Ballot States
These final ballot states allow us to quantify the impact of various risks on an election. Note that the preferred ballot status is for ballots to reach the (C,U) state.
@@ 3.3 Attacks Layer
The attacks layer is comprised of active vulnerabilities to the VBM process and can reflect different attacker goals, including changing the outcome of an election or eroding trust in political systems. Vulnerabilities in the VBM system are documented in attack trees by the US Election Assistance Commission [34] and have been expanded upon since the 2020 General Election [23]. The terminal nodes of these attack trees represent “access” points to additional vulnerabilities within the process. For this reason, we model each of these terminal leaf nodes of the attack trees. Attack trees, a valuable tool in cybersecurity, help assess threats to a system [27]. They depict potential vulnerabilities and attacker paths in a Boolean logic tree structure, where the root node represents the ultimate goal of an attacker. Attack trees use a combination of Boolean AND($\times$) and OR($+$) logic gates to trace paths from terminal attack nodes back to the root node. Figure 1 shows an example of an attack tree that visualizes two attacks corresponding to the malicious loss ($X{13}$) and accidental loss ($X{14}$) of a ballot. In Figure 1, terminal attack nodes $X{13}$ and $X{14}$ form part of an attack path leading to a successful insider attack, represented by root node 1 with only OR-gates along the paths. These attack paths offer a static view of threats within the VBM process, with the terminal leaf nodes $X{13}$ and $X{14}$ serving as entry points.
@@ 3.4 Mitigations Layer
Mitigation name Label Mitigation description Controlling entity
Automatic ballot notifications M3 The ability of a voter to attain the status of their ballot. This has a high probability if automatic notifications are provided via BalloTrax/Ballot Scout. Election office
Replacement ballots M4 Replacement ballot package request Voter
Automatic ballot reminders M5 Notify voter to send ballot back before deadline Election office
Early voting M6 In-person absentee voting Voter
Drop boxes M7 Return ballot via drop box Voter
This framework necessitates a thorough understanding of the connections between terminal node attacks in the VBM process and the available mitigation strategies designed to counteract them. To facilitate this, we employ a framework that examines the impact of terminal node attacks on the VBM process and incorporates recourse actions available to counter these attacks. Table 3 illustrates the relationship between attacks and mitigations. In this table, an “O” indicates that a mitigation strategy can counter the threat without impacting the ballot processing time. In contrast, a “D” signifies that while mitigation can counter the threat, it introduces a delay in the VBM process. For example, the scenario of a lost ballot ($X{13}$), a type of malicious attack, causes a delay resulting from the mitigation logic that includes the time required to notify the voter about their missing ballot followed by the time it takes for the voter to decide on an appropriate recourse action, such as M4 or M6. This scenario exemplifies how Table 3 summarizes key linkages in the DTMC model introduced in the next section. Consistent with the attacks in Table 1, terminal attack nodes listed in bold are incorporated into the DTMC model and are included in the computational results.
@@ 4 DTMC Model
Figure 2: Layered Network for time intervals $1\leq t<T-1$
@@ 5 Wisconsin 2020 General Election Case Study
We conduct a case study based on data from the 2020 General Election in Milwaukee County, Wisconsin. In 2020 Milwaukee County, the largest county in Wisconsin, had 939,489 residents and 478 voting wards [33, 4, 15]. Milwaukee County election officials used multiple absentee voting mechanisms during the 2020 General Election, including VBM with designated drop boxes and in-person absentee voting. The state of Wisconsin allowed absentee voting with “no excuses” and sent voter reminders to return ballots [17]. At that time, the county had 550,132 registered voters and experienced a turnout of 83.67 percent [38]. There are no known malicious attacks on VBM in Milwaukee County in the 2020 General Election, although there were two convictions elsewhere in Wisconsin associated with independent instances of VBM election fraud, each affecting a single ballot [32].
@@ 5.2 Transition probabilities
Table 4 outlines the transition probabilities for each arc active in the baseline model. The first and second columns define the two DTMC states associated with a non-zero transition probability. The next column provides a brief functional description of the ballot state transition that the arc represents. The columns labeled Interval 1, 2 and 3 represent the values of associated transition probability across time intervals 1, 2, and 3. We omit interval 4, since these transition probabilities are 0 or 111 with arcs illustrated in Figure 3. The justification column reports a reference for each transition probability value and denotes which arcs were adjusted through calibration. The arc configuration aligns with the model shown in Figure 2.
We discuss several transition probabilities. The transition probability from process state $III$ to $IV$ is 0.9380.9380.938, which reflects the proportion of election mail that is processed on-time as reported by the USPS Inspector General [20]. The rest of the ballots remain in state $III$. The transition from state $IV$ to $V$ reflects the return of completed ballots by voters (that is, the voter rate of ballot return). This transition captures the proportion of voters who return the marked ballot within a day of receiving it in the mail, which is determined as a part of the model calibration since this value is not directly recorded. MIT Election Data + Science Lab [16] provide the values for utilization of drop boxes over USPS to return ballots. Transitions from state $V$ to state $VII$, which occur with a probability of 0.5150.5150.515, represents the proportion of ballots returned by drop boxes. [16] Transitions at the end of the time interval move ballots to their final ballot states with probability 1.0. These probabilities can be gleaned from Figure 3. All other transition probabilities are zero. We calibrated the DTMC using historical data from the WEC 2020 General Election report [40]. This involved adjusting the voter rate of ballot return, the arc probability from node $IV$ to $V$, to align the output of (C,U) and (NC,U) ballots with the observed ballot return rate seen in aggregate by the state of Wisconsin. Following calibration, we validate the model’s predictive accuracy by testing it against independent data validation points for Milwaukee County. This ensures blind prediction input data such that our model fits historical data and ensures the model is capable of making accurate predictions in varying scenarios.
@@ 5.3 Calibration
Mitigation name and label Mitigation strength Justification
@@
Automatic ballot notifications (M3) 0.0265 status were not implemented
@@
Replacement ballots (M4) 0.900 replacement ballots
Automatic ballot reminders (M5) 0.740 Region used ballot return reminders
Early voting (M6) 0.400 In-person absentee was implemented
@@
Drop boxes (M7) 0.520 Elections Dataverse [3]
@@
[…]
@@ 5.4 Validation
County specific sum
v5 All sent ballots of all ballot states 2,059,130 325,547 325,351
@@ 6 Computational Results
This section discusses the baseline DTMC model results for the Milwaukee County case study. The model was coded and implemented in Matlab, and the code and data is publicly available [8]. We then study the impact of several types of malicious attacks on the established baseline. We report the expected number of ballots in each of the final ballot states—rounded to the nearest ballot—under multiple scenarios.
@@ 6.1 Baseline model
We first analyze the VBM system under normal conditions with only non-malicious attacks in the baseline model introduced in Section 5. Non-malicious attacks (see Table 4) are included in the baseline model, since they are an uncontrolled, continuous part of the VBM process. Additionally, we perform a sensitivity analysis of four of the non-malicious attacks by varying their corresponding probabilities of occurring.
@@
[…]
@@ 6.2 Malicious attack scenarios given baseline
Next, we introduce malicious attacks to the baseline. We focus on malicious attacks $X9,X13,$ and $X29$. Malicious attacks are modeled to last a single day during the 49-day time horizon to reflect feasible real-world election interference events, and we vary the attack strength of each attack to consider different scenarios. Malicious attacks therefore, modify the transition probability matrix ($P_{t^{*}}$) for a single day associated with the attack $t^{*}$. This approach allows us to assess the impact of attack timings and the mitigation strategies for countering attacks.